OpenStudy (anonymous):

Find the exact solution of the equation: |7x-3|=|7x+6| I'm not so sure how to do this. I've tried -(7x-3)=7x+6 but I don't think its right...

hartnn (hartnn):

its correct but incomplete |a| = |b| ---> |a| = b, |a| = - b which gives a=b, a=-b

hartnn (hartnn):

so, |7x-3|=|7x+6| -----> 7x-3=-(7x+6) 7x-3=7x+6

OpenStudy (anonymous):

So do I try solve those two?

hartnn (hartnn):

yes

OpenStudy (anonymous):

I've solved 7x-3=-(7x+6) which is x=-3/14 but how do I solve 7x-3=7x+6? Wouldn't 7x-7x=0?

hartnn (hartnn):

yes, which means that has no solution.

hartnn (hartnn):

and -3/14 is the only solution :)

OpenStudy (anonymous):

Oh ok, thanks a lot for your help!

Parth (parthkohli):

There are three cases in different intervals of numbers. First, both 7x - 3 and 7x + 6 are positive. That is, 7x - 3 > 0 or x > 3/7. In this interval, 7x - 3 = 7x + 6. No solution. Second, both have opposite signs. That is, 7x - 3 < 0 but 7x + 6 > 0. Or x < 3/7 but x > -6/7. -(7x - 3) = 7x + 6. -3/14 just so happens to be in this interval.

Parth (parthkohli):

If we had some other solution, it wouldn't be considered. Next time, try plugging in your answer. These questions may be tricky!

OpenStudy (anonymous):

I understand, thanks for telling me.